Acording to doctors, a healthy diet should follow the ratio for meat to vegetables of 2.5 servings to every 4 servings. If you eat 7 servings of meat a week, how many servings of vegetables should you eat?
Answer by Paul(988)

2.5/4 = x/7 -----> cross multiply
4x = 2.5(7)
4x=17.5
x = 4.375

Hence, for 7 servings you should eat 4.375 of vegetables.